### v1

# supplement

Abstract
Distributed learning offers a practical solution for the integrative analysis of multi-source datasets, especially under privacy or communication constraints. However, addressing prospective distributional heterogeneity and ensuring communication efficiency pose significant challenges on distributed statistical analysis. In this article, we focus on integrative estimation of distributed heterogeneous precision matrices, a crucial task related to joint precision matrix estimation where computation-efficient algorithms and statistical optimality theories are still underdeveloped. To tackle these challenges, we introduce a novel HEterogeneity-adjusted Aggregating and Thresholding (HEAT) approach for distributed integrative estimation. HEAT is designed to be both communication- and computation-efficient, and we demonstrate its statistical optimality by establishing the convergence rates and the corresponding minimax lower bounds under various integrative losses. To enhance the optimality of HEAT, we further propose an iterative HEAT (IteHEAT) approach. By iteratively refining the higher-order errors of HEAT estimators through multi-round communications, IteHEAT achieves geometric contraction rates of convergence. Extensive simulations and real data applications validate the numerical performance of HEAT and IteHEAT methods.
Keywords: Data Integration; Aggregating and Thresholding; Iterative Algorithm; Minimax Lower Bound.
1 Introduction
The proliferation of high-dimensional data presents numerous opportunities for statistical learning but can lead to less accurate analyses when applying targeted algorithms to a single local dataset with limited sample size. Fortunately, advancements in big data collection, storage, and transmission capabilities have made it feasible to access multiple datasets containing similar information. These datasets may originate from diverse populations, institutions and geographical locations, such as neuroimaging data (Bellec et al.,, 2017), electronic health records (Kohane et al.,, 2021), and genetic data (Zhou et al.,, 2022). By leveraging the inherent similarities among these distributed data, integrative analysis holds significant promise for enhancing statistical accuracy.
Nevertheless, synthesizing information from multiple sources is highly challenging. First, the prospective between-study heterogeneity complicates data integration. Such heterogeneity may manifest in various forms, including variations in genotype-phenotype associations (Frazier-Wood et al.,, 2013) and the differences in genetic variants across diverse ethnic populations (Gurdasani et al.,, 2019). Additionally, variability arising from inconsistent data collection and measurement across institutions further contributes to data heterogeneity (Kohane et al.,, 2021). Second, due to privacy or communication concerns, individual raw data often cannot be shared across different locations. Instead, only summary or intermediate statistics can be communicated among different sources. This constraint on data sharing may negatively impact the learning accuracy, exacerbating the challenges of data integration.
In this article, we focus on the distributed integrative estimation of high-dimensional precision matrices by taking into account the aforementioned practical challenges. The precision matrix, defined as the inverse of the covariance matrix, is central to various statistical models and plays a pivotal role in mathematical statistics. Moreover, the sparsity pattern of a precision matrix serves as a valuable tool for characterizing conditional associations between covariates, and is therefore widely employed in many scientific applications such as brain connectivity (Fransson and Marrelec,, 2008) and genetic network analysis (de la Fuente et al.,, 2004). Additionally, integrative analysis of the multi-source datasets significantly enhances the association study, while the patterns of heterogeneity facilitate differential analysis (Min et al.,, 2012; Bellec et al.,, 2017; Shojaie,, 2021). Consequently, there is a critical need to focus on the distributed setup and develop communication-efficient algorithms for estimating heterogeneous precision matrices with solid statistical guarantees.
1.1 Related Works
In recent decades, significant progress has been made in the field of high-dimensional precision matrix estimation under sparsity, including penalized maximum likelihood procedures (Yuan and Lin,, 2007; Rothman et al.,, 2008; Lam and Fan,, 2009; Ravikumar et al.,, 2011), penalized linear regression approaches (Meinshausen and Bühlmann,, 2006; Sun and Zhang,, 2013; Ren et al.,, 2015), and regularized estimating equation methods (Yuan,, 2010; Cai et al.,, 2011; Cai et al., 2016b, ). The non-negligible bias induced by the regularization in these methods has prompted numerous debiasing strategies aimed at statistical inference on the entries of precision matrices (Liu,, 2013; Janková and van de Geer,, 2015, 2017; Ning and Liu,, 2017; Neykov et al.,, 2018; Chang et al.,, 2018).
To enhance the estimation accuracy of the precision matrix within a single dataset, a rich amount of joint estimation proposals have emerged. Notable examples include $\ell_{1}/\ell_{q}$ penalized procedures, e.g., $q=2$ (Danaher et al.,, 2014; Ren et al.,, 2019) or $q=\infty$ (Honorio and Samaras,, 2010; Cai et al., 2016a, ), non-convex penalized procedures (Guo et al.,, 2011; Chun et al.,, 2015), and decomposition-based methods that assume a structure involving a common matrix and individual matrices (Hara and Washio,, 2013; Lee and Liu,, 2015); see a thorough survey in Tsai et al., (2022) and many other references therein. Nevertheless, the aforementioned joint estimation methods typically reply on solving highly non-smooth regularized optimization problems, leading to computationally intensive procedures. Additionally, due to the intrinsic complexity of precision matrix estimation, statistical optimality theories for these methods remain unknown.
In a similar context to joint estimation, there is a substantial body of literature focusing on distributed analysis under distributional homogeneity. Examples include the one-shot averaging methods (Zhang et al.,, 2015; Lee et al.,, 2017; Battey et al.,, 2018; Fan et al.,, 2019; Lv and Lian,, 2022), and the surrogate loss function methods (Wang et al.,, 2017; Jordan et al.,, 2019) with applications in distributed optimization (Fan et al.,, 2023), simultaneous inference (Yu et al.,, 2022), and non-smooth analysis (Tan et al.,, 2022). These methods achieve communication-efficiency in the sense that the dimensions of the transmitted statistics are proportional to those of target parameters. In addition, Chen et al., (2019) smooths the non-smooth objective and solves it by transmitting statistics with dimensions that are quadratic in the parameter dimension; see also Wang et al., (2019); Chen et al., (2024) for further applications. On the other hand, when model distributions are heterogeneous while target parameters remain homogeneous across sites, distributed analysis can be achieved by averaging (Zhao et al.,, 2016), weighted averaging (Gu and Chen,, 2023), or surrogate estimating equations (Duan et al.,, 2022).
Despite these advancements in homogeneous settings, statistical analysis for handling distributed heterogeneous parameters remain relatively rare. For instance, Maity et al., (2022) identifies and estimates the representative global parameter across sites, while Guo et al., (2023) focuses on robust estimation of the majority parameter. Both approaches requires only local parameter estimators from the distributed sites. Liu et al., (2024) addresses heterogeneity by fusion regularization and solves the regression problem through a first-order distributed optimization algorithm. Cai et al., (2022) introduces a quadratic surrogate loss function and transmits the first-order gradients and second-order Hessians to jointly estimate the distributed regression coefficients, with further extensions to multiple testing (Liu et al.,, 2021) and transfer learning (Li et al.,, 2023). Nonetheless, developing communication-efficient and statistically optimal methods for integrative analysis in the presence of heterogeneity remains an active area of research.
1.2 Contributions
In this article, we aim to develop distributed integrative estimation algorithms for heterogeneous precision matrices, ensuring both communication and computation efficiency along with statistically provable optimality. A novel method, HEterogeneity-adjusted Aggregating and Thresholding (HEAT), is proposed to achieve these goals effectively. Methodologically, the distributed HEAT algorithm consists of two main steps. First, an individual local estimation is derived at each local site, and summary statistics are then transmitted to the central site. Second, without any individual-level information, an integration step is conducted at the central site by performing HEAT. Theoretically, we drive the estimation rates of convergence for HEAT over a class of heterogeneous precision matrices under various norms of the introduced integrative loss matrices. Moreover, we demonstrate the statistical optimality of HEAT by establishing the matched minimax lower bounds for integrative estimation. Finally, to enhance the optimality, we further propose an iterative HEAT (IteHEAT) method that iteratively refines the higher-order errors of HEAT through multi-round communications between local sites and the central site. A contraction theory is derived for IteHEAT, demonstrating the geometric vanishing rates of higher-order errors and thereby indicating improved estimation accuracy.
The newly established methodologies and theories in this article make several valuable contributions to both distributed statistical learning and integrative precision matrix analysis. First, the proposed HEAT and IteHEAT methods introduce novel methodological advances in distributed algorithms under heterogeneity. In comparison to existing communication-efficient algorithms developed under homogeneity, our approaches demonstrate flexibility and applicability across distributional heterogeneity scenarios. Additionally, our methods provide communication advantages over distributed algorithms designed for heterogeneous settings, such as Cai et al., (2022). Second, the proposed IteHEAT algorithm provides insights into the statistical refinement of higher-order errors and demonstrates potential for improving estimation accuracy in integrative analysis. Third, the proposed methods complement the field of multi-task learning for precision matrices. Unlike the typical joint estimation approaches that rely on solving highly non-smooth optimizations, HEAT and IteHEAT are optimization-free for integration, ensuring computational efficiency. Fourth, we establish the statistical optimality theories of HEAT and IteHEAT by deriving the rates of convergence and the corresponding minimax lower bounds. To the best of our knowledge, the current minimax theories are the first among the extensive literature on joint precision matrix estimation. The introduced lower bound techniques and constructions of least favorable matrices are of independent interest and make valuable contributions to the field.
1.3 Organization
The rest of the article is organized as follows. Section 2 introduces the problem framework. Section 3 proposes the HEAT method, including the individual estimation step at the local sites and the integration step at the central site. The theoretical results of HEAT are established in Section 4.1, and the minimax lower bounds are derived in Section 4.2. Section 5 introduces the refined IteHEAT method and provides the corresponding theoretical analysis. The numerical performance of the proposed methods is evaluated through simulations in Section 6, followed by a real data analysis in Section 7. Section 8 concludes the article with discussions. All technical proofs and additional results are collected in the Supplementary Material.
2 Notations and Problem Setup
2.1 Notations
For an integer $k>0$, let $[k]=\{1,2,\ldots,k\}$; denote by $0_{k}$ and $1_{k}$ the vectors of zeros and ones with length $k$, respectively. For a set $\mathcal{A}$, define $|\mathcal{A}|$ to be its cardinality. For a vector $a\in\mathbb{R}^{n}$ and a subset $\mathcal{A}\subset[n]$, denote by $a_{\mathcal{A}}=\left(a_{i}\right)_{i\in\mathcal{A}}\in\mathbb{R}^{|\mathcal{A}|}$ a subvector of $a$. Let $\mathcal{A}^{c}=[n]\setminus\mathcal{A}$, and define $a_{-\mathcal{A}}=a_{\mathcal{A}^{c}}\in\mathbb{R}^{n-|\mathcal{A}|}$, and $a_{-j}=a_{-\{j\}}$. For a matrix $A=\left(A_{i,j}\right)_{i\in[n_{1}],j\in[n_{2}]}\in\mathbb{R}^{n_{1}\times n_{2}}$ and subsets $\mathcal{A}_{1}\subset[n_{1}],~{}\mathcal{A}_{2}\subset[n_{2}]$, let $A_{\mathcal{A}_{1},\mathcal{A}_{2}}=\left(A_{i,j}\right)_{i\in\mathcal{A}_{1},j\in\mathcal{A}_{2}}\in\mathbb{R}^{|\mathcal{A}_{1}|\times|\mathcal{A}_{2}|}$ be a submatrix of $A$. For $j\in[n_{2}]$, define $A_{\cdot,j}=A_{[n_{1}],\{j\}}$ and $A_{\cdot,-j}=A_{[n_{1}],\{j\}^{c}}$. Define $|A|_{\infty}=\max_{i\in[n_{1}],j\in[n_{2}]}|A_{i,j}|$ to be the entry-wise maximum norm, and $\|A\|_{l}=\sup_{x\in\mathbb{R}^{n_{2}}\setminus\{0_{n_{2}}\}}\|Ax\|_{l}/\|x\|_{l}$ to be the matrix $l$-norm with $l\in[1,\infty]$. Denote by $S_{+}^{n}$ the set of $n\times n$ symmetric positive definite matrices.
For two positive deterministic sequences $\{a_{n}\}$ and $\{b_{n}\}$, write $a_{n}\lesssim b_{n}$ if there exists a constant $C>0$ such that $a_{n}\leq Cb_{n}$ for all sufficiently large $n$, and $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. For two sequences of positive random variables $\{X_{n}\}$ and $\{Y_{n}\}$, write $X_{n}\lesssim_{\mathsf{P}}Y_{n}$ if $X_{n}\lesssim Y_{n}$ holds with probability tending to 1 as $n\rightarrow\infty$. For a random variable $X$, define $\|X\|_{e_{k}}=\inf\{t\geq 0:\mathbb{E}e_{k}(|X|/t)\leq 1\}$, where $e_{k}(x)=\exp(x^{k})-1$ for $k\geq 1$; $\|X\|_{e_{2}}$ ($\|X\|_{e_{1}}$) is called the sub-Gaussian (sub-exponential) norm of $X$, and $X$ is sub-Gaussian (sub-exponential) if $\|X\|_{e_{2}}$ ($\|X\|_{e_{1}}$) is finite. A random vector $X\in\mathbb{R}^{n}$ is sub-Gaussian if $a^{\scriptscriptstyle\sf T}X$ is sub-Gaussian for any deterministic vector $a\in\mathbb{R}^{n}$, and its sub-Gaussian norm is defined as $\|X\|_{e_{2}}:=\sup_{a\in\mathbb{R}^{n}:\|a\|_{2}=1}\|a^{\scriptscriptstyle\sf T}X\|_{e_{2}}$.
Let $M$ be the total number of distributed datasets. Define $a^{(\cdot)}=\left(a^{(1)},\ldots,a^{(M)}\right)^{\scriptscriptstyle\sf T}\in\mathbb{R}^{M}$ for $a^{(1)},\ldots,a^{(M)}\in\mathbb{R}$, and $a^{(\cdot,t)}=\left(a^{(1,t)},\ldots,a^{(M,t)}\right)^{\scriptscriptstyle\sf T}\in\mathbb{R}^{M}$ for $a^{(1,t)},\ldots,a^{(M,t)}\in\mathbb{R}$. For matrices $A^{(1)},\ldots,A^{(M)}\in\mathbb{R}^{n_{1}\times n_{2}}$, define $A^{(\cdot)}=\left(A^{(1)},\ldots,A^{(M)}\right)\in(\mathbb{R}^{n_{1}\times n_{2}})^{\otimes M}$; for matrices $A^{(1,t)},\ldots,A^{(M,t)}\in\mathbb{R}^{n_{1}\times n_{2}}$, define $A^{(\cdot,t)}=\left(A^{(1,t)},\ldots,A^{(M,t)}\right)\in(\mathbb{R}^{n_{1}\times n_{2}})^{\otimes M}$. Let $C,C_{0},C_{1},c,c_{0},c_{1},\ldots$ be universal positive constants that may differ from place to place.
2.2 Problem Setup
Recall that, we focus on the distributed scenarios where multiple datasets are stored separately in $M$ different sites. Denote those local datasets by $\{X^{(m)}\}_{m\in[M]}$, where $X^{(m)}=(X_{1}^{(m)},\ldots,X_{n_{m}}^{(m)})^{\scriptscriptstyle\sf T}\in\mathbb{R}^{n_{m}\times p}$ and $X_{i}^{(m)}\in\mathbb{R}^{p},i\in[n_{m}]$ are independently and identically distributed (i.i.d.) with population mean $\mu^{(m)}\in\mathbb{R}^{p}$ and covariance $\Sigma^{(m)}\in S_{+}^{p}$. Let $N=\sum_{m=1}^{M}n_{m}$ be the total number of samples, and denote by $n=\min_{m\in[M]}n_{m}$ and $n_{\max}=\max_{m\in[M]}n_{m}$ the minimal and maximal individual sample sizes, respectively. Throughout, we consider the high-dimensional regime where $(n,p)\rightarrow\infty$, and assume that there exists a constant $c>0$ such that $M\lesssim p^{c}$, which is mild under high-dimensionality.
Due to potential privacy or communication constraints on data sharing, direct merging of individual raw data is typically prohibited. Instead, only summary or intermediate statistics can be communicated from local sites to a central machine, which presents significant challenges for integration and analysis. In this context, the parameters of interest in this article are the precision matrices, defined as
$\Omega^{(m)}=(\Omega_{j,k}^{(m)})_{j,k\in[p]}:=(\Sigma^{(m)})^{-1},m\in[M],$
and we focus on their estimation under sparsity. Besides, the datasets from different sources may exhibit heterogeneous distributions, implying potentially different precision matrices across local sites. To accommodate such heterogeneity, for each $k\in[p]$, we decompose the indices of the non-zero components in the $k$-th columns into two sets:
$S_{1,k}=\{j\in[p]:\Omega_{j,k}^{(1)}=\cdots=\Omega_{j,k}^{(M)}\neq 0\},$
$S_{2,k}=\{j\in[p]:\text{ not all }\Omega_{j,k}^{(1)},\cdots,\Omega_{j,k}^{(M)}\text{ are equal}\}.$
The entries in $S_{1,k}$ suggest a common non-zero structure, while those in $S_{2,k}$ indicate the heterogeneity. These implications yield the following parameter space:
$\Theta(s_{1},s_{2})=\left\{\Omega^{(\cdot)}\in(S_{+}^{p})^{\otimes M}:\max_{k\in[p]}|S_{1,k}|\leq s_{1},\max_{k\in[p]}|S_{2,k}|\leq s_{2}\right\}.$ (1)
Our goal is to develop distributed integrative estimation procedures for the precision matrices within the space (1), while ensuring communication efficiency.
It is worth highlighting that, $s_{2}=0$ corresponds to complete homogeneity $\Omega^{(1)}=\cdots=\Omega^{(M)}$, while $s_{1}=0$ indicates the group sparsity (Yuan and Lin,, 2006; Lounici et al.,, 2011) that has been widely considered in joint precision matrix estimations (e.g., Danaher et al.,, 2014; Cai et al., 2016a, ; Ren et al.,, 2019). Therefore, the parameter space (1) is flexible and can effectively model the similarity and heterogeneity. However, this flexibility, combined with the computational and statistical challenges inherent in precision matrix estimation, further amplifies the difficulty of distributed integrative analysis.
3 Methodology
In this section, we propose a distributed integrative algorithm for heterogeneous precision matrices estimation. The algorithm consists of two main steps. First, each local site computes its individual estimate and then transmits the local summary statistics to the central site. Second, the central site performs the HEterogeneity-adjusted Aggregating and Thresholding (HEAT) step to integrate the outputs from the local sites. The proposed method is outlined in Algorithm 3, and the detailed steps are presented in Sections 3.1 and 3.2.
[…]
Step 1. Individual Local Estimation:For $m\in[M]$, at the $m$-th site, do:
  1. 1.
    Lasso regression: calculate $\widehat{\gamma}_{j}^{(m)}=\arg\min_{\gamma\in\mathbb{R}^{p-1}}\left\{\frac{1}{2n_{m}}\|\widecheck{X}_{\cdot,j}^{(m)}-\widecheck{X}_{\cdot,-j}^{(m)}\gamma\|_{2}^{2}+\lambda_{j}^{(m)}\left\|\gamma\right\|_{1}\right\}$, for $j\in[p]$, where $\widecheck{X}_{i,j}^{(m)}:=X_{i,j}^{(m)}-\widebar{X}_{j}^{(m)}$ and $\widebar{X}^{(m)}=\frac{1}{n_{m}}\sum_{i=1}^{n_{m}}X_{i}^{(m)}$.
  2. 2.
    Regularized estimation: obtain $\widehat{\Omega}^{(m)}=(\widehat{\Omega}_{j,k}^{(m)})_{j,k\in[p]}$ via (4)-(5) based on $(\widehat{\gamma}_{j}^{(m)})_{j\in[p]}$.
  3. 3.
    Debiasing: compute $\widebar{\Omega}^{(m)}=\widehat{\Omega}^{(m)}+(\widehat{\Omega}^{(m)})^{\scriptscriptstyle\sf T}-(\widehat{\Omega}^{(m)})^{\scriptscriptstyle\sf T}\widehat{\Sigma}^{(m)}\widehat{\Omega}^{(m)}$, where $\widehat{\Sigma}^{(m)}=n^{-1}_{m}\sum_{i=1}^{n_{m}}(X_{i}^{(m)}-\widebar{X}^{(m)})(X_{i}^{(m)}-\widebar{X}^{(m)})^{\scriptscriptstyle\sf T}$.
  4. 4.
    Variance estimation: compute $\widehat{v}_{j,k}^{(m)}$ via (9) for $j,k\in[p]$.
  5. 5.
    Transmit the quantities $n_{m}$, $\widebar{\Omega}^{(m)}$ and $(\widehat{v}_{j,k}^{(m)})_{j,k\in[p]}$ to the central site.
Step 2. Heterogeneity-Adjusted Aggregating and Thresholding:At the central site, do:
  1. 1.
    Integrative estimation via HEAT:
    $\widetilde{\Omega}^{(m)}=\widehat{\Gamma}+\widehat{\Lambda}^{(m)}=(\widehat{\Gamma}_{j,k})_{j,k\in[p]}+(\widehat{\Lambda}_{j,k}^{(m)})_{j,k\in[p]},~{}m\in[M],$
    with $\widehat{\Gamma}_{j,k}=T_{1,\lambda_{1,j,k}}(\widebar{\Omega}_{j,k}),~{}\widehat{\Lambda}_{j,k}^{(\cdot)}=T_{2,\lambda_{2,j,k}}(\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}),$ and $\widebar{\Omega}_{j,k}=\frac{1}{N}\sum_{m=1}^{M}n_{m}\widebar{\Omega}_{j,k}^{(m)}$, where $T_{1,\lambda}:\mathbb{R}\rightarrow\mathbb{R}$, $T_{2,\lambda}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ are thresholding functions, and $\lambda_{1,j,k},\lambda_{2,j,k}\geq 0$ are shrinkage levels.
  2. 2.
    Send $\widetilde{\Omega}^{(m)}$ back to the $m$-th local site for $m\in[M]$.
3.1 Individual Local Estimation
In the first step, we aim to obtain an individual estimation for $\Omega^{(m)}$ based on the data $X^{(m)}$ at the $m$-th local site, for each $m\in[M]$. Define the projection coefficients by
$\gamma_{j}^{(m)}=\arg\min_{\gamma\in\mathbb{R}^{p-1}}\text{Var}\left(X_{i,j}^{(m)}-X_{i,-j}^{(m)}\gamma\right),~{}j\in[p],$ (2)
and elemental calculation yields that $\gamma_{j}^{(m)}=-\Omega_{-j,j}^{(m)}/\Omega_{j,j}^{(m)}$ and $\text{Var}\left(X_{i,j}^{(m)}-X_{i,-j}^{(m)}\gamma_{j}^{(m)}\right)=1/\Omega_{j,j}^{(m)}$. Therefore, the estimation of $\Omega^{(m)}$ can be approached through the estimation of $\gamma_{j}^{(m)}$’s. Specifically, let $\widebar{X}^{(m)}=n_{m}^{-1}\sum_{i=1}^{n_{m}}X_{i}^{(m)}$ be the empirical estimation of the mean $\mu^{(m)}$, and denote by $\widecheck{X}^{(m)}$ the centralized data, i.e., $\widecheck{X}^{(m)}=\left(\widecheck{X}_{i,j}^{(m)}\right)_{i\in[n_{m}],j\in[p]}\in\mathbb{R}^{n_{m}\times p}$ with $\widecheck{X}_{i,j}^{(m)}:=X_{i,j}^{(m)}-\widebar{X}_{j}^{(m)}$. The coefficient vector $\gamma_{j}^{(m)}$ can be estimated via the node-wise Lasso procedure (Tibshirani,, 1996; Meinshausen and Bühlmann,, 2006):
$\widehat{\gamma}_{j}^{(m)}=\arg\min_{\gamma\in\mathbb{R}^{p-1}}\left\{\frac{1}{2n_{m}}\|\widecheck{X}_{\cdot,j}^{(m)}-\widecheck{X}_{\cdot,-j}^{(m)}\gamma\|_{2}^{2}+\lambda_{j}^{(m)}\left\|\gamma\right\|_{1}\right\},~{}j\in[p].$ (3)
Then the diagonal entry $\Omega_{j,j}^{(m)}$ can be estimated by $\widehat{\Omega}_{j,j}^{(m)}$ (van de Geer et al.,, 2014) through
$\widehat{\Omega}_{j,j}^{(m)}=\frac{n_{m}}{(\widecheck{X}_{\cdot,j}^{(m)})^{\scriptscriptstyle\sf T}(\widecheck{X}_{\cdot,j}^{(m)}-\widecheck{X}_{\cdot,-j}^{(m)}\widehat{\gamma}_{j}^{(m)})},$ (4)
and the off-diagonal vector $\Omega_{-j,j}^{(m)}$ can be estimated by
$\widehat{\Omega}_{-j,j}^{(m)}=-\widehat{\gamma}_{j}^{(m)}\widehat{\Omega}_{j,j}^{(m)}.$ (5)
Due to the biases induced by the regularization in (3), debiased methods have been developed to make inference on the entries of $\Omega^{(m)}$ (e.g., Liu,, 2013; Janková and van de Geer,, 2015, 2017; Ning and Liu,, 2017; Neykov et al.,, 2018). In general, the debiased estimator $\widebar{\Omega}^{(m)}$ can be represented as follows:
$\widebar{\Omega}^{(m)}=\widehat{\Omega}^{(m)}+(\widehat{\Omega}^{(m)})^{\scriptscriptstyle\sf T}-(\widehat{\Omega}^{(m)})^{\scriptscriptstyle\sf T}\widehat{\Sigma}^{(m)}\widehat{\Omega}^{(m)},$ (6)
where $\widehat{\Sigma}^{(m)}={n_{m}^{-1}}\sum_{i=1}^{n_{m}}(X_{i}^{(m)}-\widebar{X}^{(m)})(X_{i}^{(m)}-\widebar{X}^{(m)})^{\scriptscriptstyle\sf T}$ is the sample covariance. It can be shown that $\widebar{\Omega}_{j,k}^{(m)},j,k\in[p]$ is asymptotically linear with influence functions
$\phi_{i,j,k}^{(m)}:=\Omega^{(m)}_{j,k}-\Omega_{j,j}^{(m)}\Omega_{k,k}^{(m)}\epsilon_{i,j}^{(m)}\epsilon_{i,k}^{(m)},~{}i\in[n_{m}],$ (7)
where $\epsilon_{i,j}^{(m)}$ is the projection residual in (2), i.e.,
$\epsilon_{i,j}^{(m)}:=X_{i,j}^{(m)}-\mu_{j}^{(m)}-(X_{i,-j}^{(m)}-(\mu_{-j}^{(m)})^{\scriptscriptstyle\sf T})\gamma_{j}^{(m)}.$ (8)
Then the asymptotic variance of $\sqrt{n_{m}}(\widebar{\Omega}_{j,k}^{(m)}-\Omega_{j,k}^{(m)})$ is given by $v_{j,k}^{(m)}:=\text{Var}(\phi_{i,j,k}^{(m)})=\mathbb{E}(\phi_{i,j,k}^{(m)})^{2}$.
To make valid inference for $\Omega_{j,k}^{(m)}$, the variance term $v_{j,k}^{(m)}$ should be estimated consistently. In the special case where the data distribution is jointly Gaussian, elemental calculation yields that $v_{j,k}^{(m)}=(\Omega_{j,k}^{(m)})^{2}+\Omega_{j,j}^{(m)}\Omega_{k,k}^{(m)}$, and hence plug-in estimator using either $\widehat{\Omega}^{(m)}$ or $\widebar{\Omega}^{(m)}$ can be directly employed to estimate $v_{j,k}^{(m)}$. In this article, to accommodate more general distributions beyond Gaussianity, we estimate $v_{j,k}^{(m)}$ through the empirical variance of the estimated influence functions:
$\widehat{v}_{j,k}^{(m)}=\frac{1}{n_{m}}\sum_{i=1}^{n_{m}}(\widehat{\phi}_{i,j,k}^{(m)})^{2},$ (9)
where $\widehat{\phi}_{i,j,k}^{(m)}$ is derived via
$\widehat{\phi}_{i,j,k}^{(m)}=\widebar{\Omega}_{j,k}^{(m)}-\widebar{\Omega}_{j,j}^{(m)}\widebar{\Omega}_{k,k}^{(m)}\widehat{\epsilon}_{i,j}^{(m)}\widehat{\epsilon}_{i,k}^{(m)},\text{ and }\widehat{\epsilon}_{i,j}^{(m)}=\widecheck{X}_{i,j}^{(m)}-\widecheck{X}_{i,-j}^{(m)}\widehat{\gamma}_{j}^{(m)}.$ (10)
Finally, the $m$-th site transmits the sample size $n_{m}$, the local estimator $\widebar{\Omega}^{(m)}$ in (6), and the variance estimator $(\widehat{v}_{j,k}^{(m)})_{j,k\in[p]}$ in (9) to the central site, followed by the integration step in Section 3.2.
3.2 Heterogeneity-Adjusted Aggregating and Thresholding
To exploit the data similarity and improve the estimation accuracy, we next perform a heterogeneity-adjusted aggregating and thresholding step based on the local estimation in the previous section.
We start with a reparametrization of the matrix $\Omega^{(m)}$ by decomposing it into two components: a common baseline matrix that indicates the between-sites similarity and an individual matrix that accommodates the heterogeneity. To be specific, we write
$\Omega^{(m)}=\Gamma+\Lambda^{(m)},~{}m\in[M],$ (11)
where $\Gamma=(\Gamma_{j,k})_{j,k\in[p]}$ is the common baseline matrix, $\Lambda^{(m)}=({\Lambda}_{j,k}^{(m)})_{j,k\in[p]}$ reflects distribution heterogeneity, and we impose additional equality constraints $\sum_{m=1}^{M}n_{m}\Lambda_{j,k}^{(m)}=0,j,k\in[p]$ for identification. Such representation is common in the analysis of variance. It follows that $\Gamma=\frac{1}{N}\sum_{m=1}^{M}n_{m}\Omega^{(m)}$ is the weighted average of $\Omega^{(m)},m\in[M]$, and that $\Lambda_{j,k}^{(1)}=\cdots=\Lambda_{j,k}^{(M)}$ implies $\Lambda_{j,k}^{(\cdot)}=0_{M}$ and $\Omega_{j,k}^{(1)}=\cdots=\Omega_{j,k}^{(M)}=\Gamma_{j,k}$. Consequently, the heterogeneous patterns in the space (1) are characterized by $\Lambda^{(\cdot)}$.
Based on the heterogeneity representation in (11), we propose the HEAT approach to integrate the results from local sites. Specifically, the HEAT estimators for $\Omega^{(m)},m\in[M]$ are calculated by
$\widetilde{\Omega}^{(m)}=\widehat{\Gamma}+\widehat{\Lambda}^{(m)},m\in[M],$ (12)
where the $\widehat{\Gamma}=(\widehat{\Gamma}_{j,k})_{j,k\in[p]}$ and $\widehat{\Lambda}^{(m)}=(\widehat{\Lambda}_{j,k}^{(m)})_{j,k\in[p]}$ are given by
$\widehat{\Gamma}_{j,k}=T_{1,\lambda_{1,j,k}}(\widebar{\Omega}_{j,k}),$ (13)
$\widehat{\Lambda}_{j,k}^{(\cdot)}=T_{2,\lambda_{2,j,k}}(\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}),$ (14)
with $\widebar{\Omega}_{j,k}=\frac{1}{N}\sum_{m=1}^{M}n_{m}\widebar{\Omega}_{j,k}^{(m)}$. Here $T_{1,\lambda}:\mathbb{R}\rightarrow\mathbb{R}$ and $T_{2,\lambda}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ are some thresholding functions, and $\lambda_{1,j,k},\lambda_{2,j,k}\geq 0$ are certain shrinkage levels. Finally, the HEAT estimator $\widetilde{\Omega}^{(m)}$ is transmitted back to the $m$-th local site for all $m\in[M]$.
We now briefly explain the proposed HEAT method above. Step (12) can essentially be viewed as a double shrinkage approach for the individual estimators. By the representation (11), smaller absolute values of the components in $\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}$ indicate stronger signals of homogeneity. Therefore, shrinkage towards zero is performed by $T_{2,\lambda_{2,j,k}}$ to adjust for heterogeneity. In the meanwhile, the thresholding function $T_{1,\lambda_{1,j,k}}$ shrinks the estimator $\widebar{\Omega}_{j,k}$ to encourage the sparsity patterns of the common matrix $\Gamma$.
Next, we discuss the choices of the thresholding functions and the shrinkage levels. By Rothman et al., (2009), a univariate thresholding function $T_{1,\lambda}$ should satisfy the following properties:
$T_{1,\lambda}(x)=0\text{ if }|x|\leq\lambda,\text{ and }\left|T_{1,\lambda}(x)-x\right|\leq\lambda\text{ for all }x\in\mathbb{R}.$ (15)
Proper functions include soft, hard, SCAD (Fan and Li,, 2001) and MCP (Zhang,, 2010) thresholding. For the multivariate $T_{2,\lambda}$, we generalize the univariate criteria (15) and propose the following multivariate thresholding rules:
$T_{2,\lambda}(x)=0_{M}\text{ if }\|x\|\leq\lambda,\text{ and }\left\|T_{2,\lambda}(x)-x\right\|\leq\lambda\text{ for all }x\in\mathbb{R}^{M},$ (16)
[…]
For the shrinkage levels $\lambda_{1,j,k}$ and $\lambda_{2,j,k}$ in (13)-(14), both universal shrinkage (e.g., Bickel and Levina,, 2008; Rothman et al.,, 2009; Cai and Zhou,, 2012) and entry-dependent adaptive shrinkage (e.g., Cai and Liu,, 2011; Fan et al.,, 2013) can be applied. In this article, we employ the entry-dependent levels to accommodate the entry-wise volatility heteroscedasticity for $\widebar{\Omega}_{j,k}$ and $\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}$, as shown in Propositions 1-2 of Section 4.1.
Remark 1.
The HEAT method achieves both communication efficiency in distributed settings and computation efficiency for joint precision matrix estimation. On the one hand, the dimensions of the transmitted summaries in Algorithm 3 are proportional to the dimension of the target parameters. It aligns with the existing communication-efficient algorithms developed under homogeneity (e.g., Lee et al.,, 2017; Battey et al.,, 2018; Jordan et al.,, 2019) and shows advantages over the heterogeneous approach in Cai et al., (2022), where $p\times p$ matrices are transmitted even though the target parameters are $p$-dimensional. On the other hand, unlike existing joint precision matrix estimation algorithms that involve highly non-smooth optimizations and computationally intensive procedures, HEAT is optimization-free for integration, and therefore is computationally efficient.
4 Theoretical Properties
This section presents the statistical convergence rates for Algorithm 3 and shows its optimality by establishing the minimax information limits for the current integrative estimation task. To begin with, we introduce proper measures to evaluate the estimation accuracy. For $A^{(\cdot)}=(A^{(1)},\cdots,A^{(M)})\in(\mathbb{R}^{p\times p})^{\otimes M}$ and $B^{(\cdot)}=(B^{(1)},\cdots,B^{(M)})\in(\mathbb{R}^{p\times p})^{\otimes M}$, define the integrative loss matrices $L_{1,r}(A^{(\cdot)},B^{(\cdot)})\in\mathbb{R}^{p\times p}$ and $L_{2,r}(A^{(\cdot)},B^{(\cdot)})\in\mathbb{R}^{p\times p}$ as
$\left(L_{1,r}(A^{(\cdot)},B^{(\cdot)})\right)_{j,k}=\|A_{j,k}^{(\cdot)}-B_{j,k}^{(\cdot)}\|_{1,w}^{r},~{}j,k\in[p],$ (17)
$\left(L_{2,r}(A^{(\cdot)},B^{(\cdot)})\right)_{j,k}=\|A_{j,k}^{(\cdot)}-B_{j,k}^{(\cdot)}\|_{2,w}^{r},~{}j,k\in[p],$ (18)
where $r\geq 1$ is a fixed constant, and the weighted $\ell_{1}$ norm $\|\cdot\|_{1,w}$ and weighted $\ell_{2}$ norm $\|\cdot\|_{2,w}$ for a vector $a\in\mathbb{R}^{M}$ are given by $\|a\|_{1,w}:=\frac{1}{N}\sum_{m=1}^{M}n_{m}|a_{m}|$ and $\|a\|_{2,w}:=\sqrt{\frac{1}{N}\sum_{m=1}^{M}n_{m}a_{m}^{2}}$. Therefore, (17)-(18) quantify the entry-wise integrative deviation between $A^{(\cdot)}$ and $B^{(\cdot)}$ under the $r$-th power of the weighted $\ell_{1}$ and $\ell_{2}$ norms, respectively. It is noteworthy that the sample size heterogeneity across the local sites is fully accommodated by the weights $\{n_{m}/N\}_{m\in[M]}$, and consequently there is no necessity to impose the commonly assumed balanced sample size condition to derive the convergence rates.
4.1 Rates of Convergence for HEAT
We first introduce a few technical assumptions.
Assumption 1.
There exist constants $C_{\max},C_{\min}>0$ such that $\min_{m\in[M]}\sigma_{\min}\left(\Sigma^{(m)}\right)\geq C_{\min}$ and $\max_{m\in[M]}\max_{j\in[p]}\Sigma_{j,j}^{(m)}\leq C_{\max}$, where $\sigma_{\min}\left(\Sigma^{(m)}\right)$ denotes the minimal eigenvalue of $\Sigma^{(m)}$.
Assumption 2.
$X_{i,j}^{(m)}-\mu_{j}^{(m)}$ and $\epsilon_{i,j}^{(m)}$ are sub-Gaussian for all $j\in[p]$ and $m\in[M]$ with uniformly bounded sub-Gaussian norms.
Assumption 3.
There exists a constant $c>0$ such that $v_{j,k}^{(m)}:=\text{Var}(\phi_{i,j,k}^{(m)})=\mathbb{E}(\phi_{i,j,k}^{(m)})^{2}\geq c$ uniformly for all $j,k\in[p]$ and $m\in[M]$, where $\phi_{i,j,k}^{(m)}$ is defined in (7).
Assumption 4.
The thresholding function $T_{1,\lambda}:\mathbb{R}\rightarrow\mathbb{R}$ satisfies the properties (15), and the function $T_{2,\lambda}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ satisfies the properties (16) with $\|\cdot\|=\|\cdot\|_{2,w}$ on $\mathbb{R}^{M}$.
Remark 2.
Assumptions 1-3 impose mild conditions on data distributions, and Assumption 4 summarizes the requirements for the thresholding functions. Specifically, Assumption 1 is a regularity condition to prevent singular covariances and avoid diverging variances. Assumption 2 excludes the heavy-tailed distributions, and Assumption 3 assumes non-degenerate variances for the influence functions $\phi_{i,j,k}^{(m)}$’s. Assumptions 2-3 can be implied by Assumption 1 if the distributions are Gaussian. The weighted norm $\|\cdot\|_{2,w}$ for $T_{2,\lambda}$ in Assumption 4 accommodates the sample size differences among the distributed sites. Section LABEL:sec:choice_threshold_fun of the supplement presents specific examples of $T_{1,\lambda}$ and $T_{2,\lambda}$ that satisfy Assumption 4, including soft, hard, SCAD and MCP thresholding functions.
Next we establish the theoretical results for the HEAT approach under the integrative losses (17)-(18) with various matrix norms. We begin with several propositions to provide useful guidance on the determination of the shrinkage levels $\lambda_{1,j,k}$ and $\lambda_{2,j,k}$ in the thresholding functions (13)-(14). The key is to provide explicit and tight quantifications for the uncertainties of the aggregated statistics $\widebar{\Omega}_{j,k}$ and $\|\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}\|_{2,w}$. Recall that $s_{1}$ and $s_{2}$ are the sparsity levels in the parameter space (1), and let $s_{0}:=s_{1}\vee s_{2}$.
Proposition 1.
Under Assumptions 1-3, assume $s_{0}\log(p)=o(n)$, $\log^{3}(p)=o(N)$, $n\geq Cs_{0}^{2}\log(p)$ for some constant $C>0$, and $\lambda_{j}^{(m)}\asymp\sqrt{\log(p)/n_{m}}$ uniformly for $j\in[p]$ and $m\in[M]$, and define
$\widebar{\lambda}_{1,j,k}=2\sqrt{\|v_{j,k}^{(\cdot)}\|_{1,w}\frac{\log(p)}{N}}+\widebar{C}_{1}\frac{s_{0}M\log(p)}{N}$
for some constant $\widebar{C}_{1}>0$, then the event $\bigcap_{j,k\in[p]}\left\{|\widebar{\Omega}_{j,k}-\Gamma_{j,k}|\leq\widebar{\lambda}_{1,j,k}\right\}$ holds with probability tending to 1.
Proposition 2.
Under Assumptions 1-2, assume $s_{0}\log(p)=o(n)$, $n\geq Cs_{0}^{2}\log(p)$ for some constant $C>0$, and $\lambda_{j}^{(m)}\asymp\sqrt{\log(p)/n_{m}}$ uniformly for $j\in[p]$ and $m\in[M]$, and define
$\widebar{\lambda}_{2,j,k}=\sqrt{\frac{\|v_{j,k}^{(\cdot)}\|_{1}+(2\sqrt{2}+\widebar{\delta})\|v_{j,k}^{(\cdot)}\|_{2}\sqrt{\log(p)}+(4+\widebar{\delta})\|v_{j,k}^{(\cdot)}\|_{\infty}\log(p)}{N}}+\widebar{C}_{2}\frac{s_{0}\sqrt{M}\log(p)}{\sqrt{nN}}+\frac{\tau}{\sqrt{N}}$
and $\tau=\widebar{C}_{3}M^{5/2}\log(p\vee n_{\max})\log(p)/n^{1/2}$ for some constants $\widebar{C}_{2},\widebar{C}_{3}>0$, then for any constant $\widebar{\delta}>0$, the event $\bigcap_{j,k\in[p]}\left\{\|\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}-\Lambda_{j,k}^{(\cdot)}\|_{2,w}\leq\widebar{\lambda}_{2,j,k}\right\}$ holds with probability tending to 1.
Propositions 1-2 characterize the volatility of $\widebar{\Omega}_{j,k}$ and $\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}$. As a result, $|\widebar{\Omega}_{j,k}|\leq\widebar{\lambda}_{1,j,k}$ and $\|\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}\|_{2,w}\leq\widebar{\lambda}_{2,j,k}$ respectively indicate that signals $\Gamma_{j,k}$ and $\Lambda_{j,k}^{(\cdot)}$ are weak, allowing for thresholding $\widebar{\Omega}_{j,k}$ or $\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}$ to exactly zero. The variance terms $\|v_{j,k}^{(\cdot)}\|_{1,w}$, $\|v_{j,k}^{(\cdot)}\|_{1},\|v_{j,k}^{(\cdot)}\|_{2},\|v_{j,k}^{(\cdot)}\|_{\infty}$ in $\widebar{\lambda}_{1,j,k}$ and $\widebar{\lambda}_{2,j,k}$ reflect entry-dependent volatility. Although these terms can be replaced by a universal constant for universal shrinkage levels, the proposed thresholds adapt to heteroscedasticity and provide practical guidance on shrinkage level selection. The remaining terms, $s_{0}M\log(p)/N$ in Proposition 1 and $s_{0}\sqrt{M}\log(p)/\sqrt{nN}$ in Proposition 2, represent higher-order errors of the local estimators (6). The additional term $\tau/\sqrt{N}$ in Proposition 2 is technical and potentially subject to further theoretical refinement. It arises from the Chi-square approximation for the sum-of-square-type quantity $\|\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}-\Lambda_{j,k}^{(\cdot)}\|_{2,w}$ as presented in Lemma LABEL:lemma:ss_chis_bound of the supplement. It is negligible if $M^{4}\log^{4}(p\vee n_{\max})=o(n)$, a simplification assumed in subsequent analyses for clarity. It’s essential to highlight that compared to typical concentration results for quadratic forms or polynomials, the Chi-square approximation developed here offers tight bounds on both constant terms and convergence rates. Therefore, it facilitates both practical implementation and theoretical error analysis.
In practice, the variances are unknown and we can replace them by the empirical estimator $\widehat{v}_{j,k}^{(m)}$ proposed in (9) of Section 3.1. The following Proposition 3 establishes the uniform rate of convergence for the variance estimation.
Proposition 3.
Under Assumptions 1-2, if $s_{0}\log(p)=o(n)$, $n\geq C\log(p)(s_{0}\vee\log(p))^{2}$ for some constant $C>0$, and $\lambda_{j}^{(m)}\asymp\sqrt{\log(p)/n_{m}}$ uniformly for $j\in[p]$ and $m\in[M]$, then the variance estimator $\widehat{v}_{j,k}^{(m)}$ in (9) satisfies
$|\widehat{v}_{j,k}^{(m)}-v_{j,k}^{(m)}|\lesssim_{\mathsf{P}}\max\left(a_{n}^{(m)},(a_{n}^{(m)})^{2}\right)$
uniformly for $j,k\in[p]$ and $m\in[M]$, where $a_{n}^{(m)}=\log(p\vee n_{\max})\sqrt{\left(s_{0}\vee\log(p)\right)/n_{m}}$.
By Propositions 1-3, for a small constant $\delta>0$, the shrinkage levels $\lambda_{1,j,k}$ and $\lambda_{2,j,k}$ in (13)-(14) can be determined by
$\lambda_{1,j,k}=(2+\delta)\sqrt{\|\widehat{v}_{j,k}^{(\cdot)}\|_{1,w}\frac{\log(p)}{N}}+\widebar{C}_{1}\frac{s_{0}M\log(p)}{N},$ (19)
$\lambda_{2,j,k}=(1+\delta)\sqrt{\frac{\|\widehat{v}_{j,k}^{(\cdot)}\|_{1}+2\sqrt{2}\|\widehat{v}_{j,k}^{(\cdot)}\|_{2}\sqrt{\log(p)}+4\|\widehat{v}_{j,k}^{(\cdot)}\|_{\infty}\log(p)}{N}}+\widebar{C}_{2}\frac{\sqrt{M}s_{0}\log(p)}{\sqrt{nN}}.$ (20)
Subsequently, the following theorems establish the rates of convergence for the HEAT estimator $\widetilde{\Omega}^{(\cdot)}$ under the loss matrices $L_{1,r}$ and $L_{2,r}$ in (17)-(18). Specifically, Theorem 1 evaluates the integrative estimation losses under the general matrix $l$-norm with $l\in[1,\infty]$, while Theorem 2 investigates the matrix Frobenius norm, which essentially vectorizes the loss matrices and measures the accuracy through vector $\ell_{2}$ norm.
Theorem 1.
Under Assumptions 1-4, if $n\geq Cs_{0}^{2}\log(p)$ for some constant $C>0$, $\log^{2}(p\vee n_{\max})(s_{0}\vee\log(p))=o(n)$, $M^{4}\log^{4}(p\vee n_{\max})=o(n)$, and $\lambda_{j}^{(m)}\asymp\sqrt{\log(p)/n_{m}}$ uniformly for $j\in[p]$ and $m\in[M]$, then for any $r\geq 1$ and $l\in[1,\infty]$, we have
$\|L_{1,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{l}\lesssim_{\mathsf{P}}s_{1}(\varphi_{1})^{r}+s_{2}(\varphi_{2})^{r},$
$\|L_{2,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{l}\lesssim_{\mathsf{P}}s_{1}(\varphi_{1})^{r}+s_{2}(\varphi_{2})^{r},$
where $\varphi_{1}=\sqrt{\frac{\log(p)}{N}}+\frac{s_{0}M\log(p)}{N}$ and $\varphi_{2}=\sqrt{\frac{M+\log(p)}{N}}+\frac{s_{0}\sqrt{M}\log(p)}{\sqrt{nN}}$.
Theorem 2.
Under the conditions in Theorem 1, we have
$\frac{1}{p}\|L_{1,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{\text{F}}^{2}\lesssim_{\mathsf{P}}s_{1}(\varphi_{1})^{2r}+s_{2}(\varphi_{2})^{2r},$
$\frac{1}{p}\|L_{2,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{\text{F}}^{2}\lesssim_{\mathsf{P}}s_{1}(\varphi_{1})^{2r}+s_{2}(\varphi_{2})^{2r}.$
The convergence rates in Theorem 1 consist of two components: $s_{1}(\varphi_{1})^{r}$ and $s_{2}(\varphi_{2})^{r}$, which respectively reflect the estimation errors under the homogeneity and the heterogeneity. The same decomposition and interpretation also apply to Theorem 2. Note that both $\varphi_{1}$ and $\varphi_{2}$ consist of two terms. The terms $\sqrt{\frac{\log(p)}{N}}$ and $\sqrt{\frac{M+\log(p)}{N}}$ are the leading statistical errors for integration, while $\frac{s_{0}M\log(p)}{N}$ and $\frac{\sqrt{M}s_{0}\log(p)}{\sqrt{nN}}$ are the higher-order errors induced by individual estimation, representing the additional costs of distributed data analysis. Consequently, the higher-order errors will be dominated by the leading error terms, i.e.,
$\varphi_{1}\asymp\psi_{1}:=\sqrt{\frac{\log(p)}{N}}\text{ and }\varphi_{2}\asymp\psi_{2}:=\sqrt{\frac{M+\log(p)}{N}},$
if $\frac{M^{2}}{N}\lesssim{(s_{0}^{2}\log(p))}^{-1}$ and $s_{0}^{2}\log^{2}(p)\lesssim(1\vee\frac{\log(p)}{M})n$ hold. These restrictions imply that the number of distributed sites can not be too large, and similar requirements also appear in the distributed algorithms under homogeneity (e.g., Zhang et al.,, 2015; Lee et al.,, 2017; Battey et al.,, 2018; Jordan et al.,, 2019) and heterogeneity (e.g., Maity et al.,, 2022; Cai et al.,, 2022). In the following section, it will be shown that $\psi_{1}$ and $\psi_{2}$ achieve the minimax optimal rates.
4.2 Minimax Lower Bounds for Integrative Estimation
In this section, the minimax lower bounds are established for the integrative estimation to demonstrate the optimality of HEAT approach under various loss criteria. Whenever there is no ambiguity, we abbreviate the parameter space $\Theta(s_{1},s_{2})$ defined in (1) to $\Theta$.
Theorem 3.
Suppose that $X_{i}^{(m)}\sim N(\mu^{(m)},\Sigma^{(m)})$ for $i\in[n_{m}],m\in[M]$ and Assumption 1 holds. If $23\leq s_{1}<p/$ 2, $22\leq s_{2}<p/$ 2, $s_{0}^{2}\log^{3}(p)=o(N)$, $s_{2}^{2}M=o(N)$, $N\lesssim nM$ and $p\gtrsim N^{\beta}$ for some $\beta>1$, then for any $r\geq 1$ and $l\in[1,\infty]$,
$\inf_{\widetilde{\Omega}^{(\cdot)}}\sup_{\Omega^{(\cdot)}\in\Theta}\mathbb{E}\|L_{1,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{l}\gtrsim s_{1}\psi_{1}^{r}+s_{2}\psi_{2}^{r},$
$\inf_{\widetilde{\Omega}^{(\cdot)}}\sup_{\Omega^{(\cdot)}\in\Theta}\mathbb{E}\|L_{2,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{l}\gtrsim s_{1}\psi_{1}^{r}+s_{2}\psi_{2}^{r}.$
Theorem 4.
Under the conditions in Theorem 3, we have
$\inf_{\widetilde{\Omega}^{(\cdot)}}\sup_{\Omega^{(\cdot)}\in\Theta}\frac{1}{p}\mathbb{E}\|L_{1,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{\text{F}}^{2}\gtrsim s_{1}\psi_{1}^{2r}+s_{2}\psi_{2}^{2r},$
$\inf_{\widetilde{\Omega}^{(\cdot)}}\sup_{\Omega^{(\cdot)}\in\Theta}\frac{1}{p}\mathbb{E}\|L_{2,r}(\widetilde{\Omega}^{(\cdot)},\Omega^{(\cdot)})\|_{\text{F}}^{2}\gtrsim s_{1}\psi_{1}^{2r}+s_{2}\psi_{2}^{2r}.$
Theorems 3-4 establish the minimax lower bounds under the matrix $l$-norm and Frobenius norm, respectively. The condition $p\gtrsim N^{\beta}$ is technical and serves to quantify the difficulty of distinguishing the least favorable parameters by calculating the divergences between carefully constructed distributions. Despite the extensive literature on integrative precision matrix estimation as reviewed in Section 1.1, to the best of our knowledge, Theorems 3-4 are among the first to establish the minimax theories. The lower bound techniques and the matrix constructions in the proofs of Theorems 3-4 are novel and are of independent interest to the integrative precision matrix analysis under heterogeneity.
By combining Theorems 3-4 and Theorems 1-2, it is shown that HEAT achieves rate-optimality under various integrative losses subject to certain restrictions on the number of distributed sites. In the following section, we will introduce a refined HEAT method which attains enhanced estimation accuracy and optimality through iterative communications between the central and local sites.
5 Iterative HEAT Estimation
As demonstrated in Section 4, achieving optimality for HEAT requires additional constraints on the number of distributed sites due to the higher-order errors of aggregated estimators. Moreover, even with a small or moderately large number of distributed sites, reducing these errors can enhance numerical performance in practice. To address this, we further propose the iterative HEAT (IteHEAT) method, which iteratively refines the higher-order errors of HEAT through multi-round communications between central and local sites. Notably, unlike the existing iterative distributed algorithms developed for homogeneous settings (e.g., Wang et al.,, 2017; Jordan et al.,, 2019; Chen et al.,, 2019), addressing distributional heterogeneity presents distinct challenges and necessitates both methodological and theoretical innovations.
In the following, we outline the proposed IteHEAT in Algorithm 5 and provide detailed descriptions and theoretical properties in Sections 5.1 and 5.2.
[…]
First-Round Communication:
  1. 1.
    Sample Splitting:
    For $m\in[M]$, at the $m$-th site, do:
    1. (a)
      Randomly split the samples into two disjoint sets $\mathcal{I}^{(m_{1})}$ and $\mathcal{I}^{(m_{2})}$ such that $[n_{m}]=\mathcal{I}^{(m_{1})}\bigcup\mathcal{I}^{(m_{2})}$ and $|\mathcal{I}^{(m_{2})}|=\kappa_{m}n_{m}$ for some constant $\kappa_{m}\in(0,1)$.
    2. (b)
      Compute $\widehat{\Omega}^{(m_{1})}$ in (21) based on $\{X_{i}^{(m)},i\in\mathcal{I}^{(m_{1})}\}$, and construct $\widehat{\Sigma}^{(m_{2})}$ based on $\{X_{i}^{(m)},i\in\mathcal{I}^{(m_{2})}\}$.
    3. (c)
      Transmit the proportion $\kappa_{m}$ to the central site.
  2. 2.
    Initialization:
    Implement Algorithm 3, and obtain the first-round HEAT estimator $\widetilde{\Omega}^{(m)}$ for $m\in[M]$. Define $\widetilde{\Omega}^{(m,1)}:=\widetilde{\Omega}^{(m)}$, $m\in[M]$.
Multi-Round Communications:For the iterations $t=2,\ldots,T$, do:
  1. 1.
    Individual Local Estimation:
    For $m\in[M]$, at the $m$-th site, do:
    1. (a)
      Symmetrized debiased estimation:
      $\widebar{\Omega}^{(m,t)}=\frac{1}{2}\left(\widebar{\Omega}^{(m,t-1/2)}+(\widebar{\Omega}^{(m,t-1/2)})^{\scriptscriptstyle\sf T}\right),$
      where $\widebar{\Omega}^{(m,t-1/2)}=(\widehat{\Omega}^{(m_{1})})^{\scriptscriptstyle\sf T}+\widetilde{\Omega}^{(m,t-1)}-(\widehat{\Omega}^{(m_{1})})^{\scriptscriptstyle\sf T}\widehat{\Sigma}^{(m_{2})}\widetilde{\Omega}^{(m,t-1)}$.
    2. (b)
      Transmit $\widebar{\Omega}^{(m,t)}$ to the central site.
  2. 2.
    IteHEAT Estimation:
    At the central site, do:
    1. (a)
      Integrative estimation via HEAT:
      $\widetilde{\Omega}^{(m,t)}=\widehat{\Gamma}^{(t)}+\widehat{\Lambda}^{(m,t)}=(\widehat{\Gamma}_{j,k}^{(t)})_{j,k\in[p]}+(\widehat{\Lambda}_{j,k}^{(m,t)})_{j,k\in[p]},$
      where
      $\widehat{\Gamma}_{j,k}^{(t)}=T_{1,\lambda_{1,j,k}^{(t)}}(\widebar{\Omega}_{j,k}^{(t)}),\widehat{\Lambda}_{j,k}^{(\cdot,t)}=T_{2,\lambda_{2,j,k}^{(t)}}(\widebar{\Omega}_{j,k}^{(\cdot,t)}-\widebar{\Omega}_{j,k}^{(t)}1_{M}),$
      and $\widebar{\Omega}^{(t)}=\frac{1}{N}\sum_{m=1}^{M}n_{m}\widebar{\Omega}^{(m,t)}$.
    2. (b)
      Send $\widetilde{\Omega}^{(m,t)}$ back to the $m$-th local site for $m\in[M]$.
5.1 IteHEAT Algorithm
IteHEAT begins with a preliminary data splitting step and initialization using Algorithm 3, and then refines the estimates through iterative local debiasing and central integration. This approach offers insights into the statistical refinement of higher-order errors and the accuracy enhancement in integrative analysis.
First, at each site, the dataset is randomly split into two disjoint subsets, and then these two sets are separately used to construct the regularized estimation for $\Omega^{(m)}$ and the covariance estimation for $\Sigma^{(m)}$. Specifically, for each $m\in[M]$, the $m$-th site randomly splits the samples into two disjoint sets, $\mathcal{I}_{1}^{(m_{1})}$ and $\mathcal{I}^{(m_{2})}$, such that $[n_{m}]=\mathcal{I}^{(m_{1})}\bigcup\mathcal{I}^{(m_{2})}$ and $|\mathcal{I}^{(m_{2})}|=\kappa_{m}n_{m}$, where the splitting proportion $\kappa_{m}\in(0,1)$ is a preassigned universal constant. The superscripts $m_{1}$ and $m_{2}$ indicate the first and second subsets at the $m$-th site, respectively. For $a\in\{1,2\}$, we define the subsets $X^{(m_{a})}:=X_{\mathcal{I}^{(m_{a})},[p]}^{(m)}\in\mathbb{R}^{|\mathcal{I}^{(m_{a})}|\times p}$ and the centralized subsets $\widecheck{X}^{(m_{a})}=(\widecheck{X}^{(m_{a})}_{i,j})\in\mathbb{R}^{|\mathcal{I}^{(m_{a})}|\times p}$, where $\widecheck{X}^{(m_{a})}_{i,j}=X^{(m_{a})}_{i,j}-\widebar{X}^{(m_{a})}_{j}$ and $\widebar{X}^{(m_{a})}=\frac{1}{|\mathcal{I}^{(m_{a})}|}\sum_{i\in\mathcal{I}^{(m_{a})}}X_{i}^{(m)}$. Similar to the local estimator $\widehat{\Omega}^{(m)}$ in (4)-(5), the regularized estimation $\widehat{\Omega}^{(m_{1})}$ is constructed based on the first subset $X^{(m_{1})}$:
$\widehat{\Omega}_{j,j}^{(m_{1})}=\frac{(1-\kappa_{m})n_{m}}{(\widecheck{X}_{\cdot,j}^{(m_{1})})^{\scriptscriptstyle\sf T}(\widecheck{X}_{\cdot,j}^{(m_{1})}-\widecheck{X}_{\cdot,-j}^{(m_{1})}\widehat{\gamma}_{j}^{(m_{1})})},\text{ and }\widehat{\Omega}_{-j,j}^{(m_{1})}=-\widehat{\gamma}_{j}^{(m_{1})}\widehat{\Omega}_{j,j}^{(m_{1})},$ (21)
where
$\widehat{\gamma}_{j}^{(m_{1})}=\arg\min_{\gamma\in\mathbb{R}^{p-1}}\left\{\frac{1}{2(1-\kappa_{m})n_{m}}\|\widecheck{X}_{\cdot,j}^{(m_{1})}-\widecheck{X}_{\cdot,-j}^{(m_{1})}\gamma\|_{2}^{2}+\frac{\lambda_{j}^{(m)}}{\sqrt{1-\kappa_{m}}}\left\|\gamma\right\|_{1}\right\}.$ (22)
Besides, the empirical covariance $\widehat{\Sigma}^{(m_{2})}:=\frac{1}{|\mathcal{I}^{(m_{2})}|}(\widecheck{X}^{(m_{2})})^{\scriptscriptstyle\sf T}\widecheck{X}^{(m_{2})}$ is constructed based on the second subsets. For $m\in[M]$, the proportion $\kappa_{m}$ is sent to the central site.
Next, Algorithm 3 is implemented to obtain the initial first-round HEAT estimator $\widetilde{\Omega}^{(\cdot,1)}:=\widetilde{\Omega}^{(\cdot)}$, followed by multi-round communications to achieve the IteHEAT estimation. Specifically, for $t\geq 2$, during the $t$-th round of communication, the $m$-th site computes the symmetrized debiased estimator by
$\widebar{\Omega}^{(m,t)}=\frac{1}{2}\left(\widebar{\Omega}^{(m,t-1/2)}+(\widebar{\Omega}^{(m,t-1/2)})^{\scriptscriptstyle\sf T}\right),$ (23)
where
$\widebar{\Omega}^{(m,t-1/2)}=(\widehat{\Omega}^{(m_{1})})^{\scriptscriptstyle\sf T}+\widetilde{\Omega}^{(m,t-1)}-(\widehat{\Omega}^{(m_{1})})^{\scriptscriptstyle\sf T}\widehat{\Sigma}^{(m_{2})}\widetilde{\Omega}^{(m,t-1)}.$ (24)
Then, $\widebar{\Omega}^{(m,t)}$ is transmitted to the central site for all $m\in[M]$, followed by the HEAT estimation:
$\widetilde{\Omega}^{(m,t)}=\widehat{\Gamma}^{(t)}+\widehat{\Lambda}^{(m,t)}=(\widehat{\Gamma}_{j,k}^{(t)})_{j,k\in[p]}+(\widehat{\Lambda}_{j,k}^{(m,t)})_{j,k\in[p]},$ (25)
where
$\widehat{\Gamma}_{j,k}^{(t)}=T_{1,\lambda_{1,j,k}^{(t)}}(\widebar{\Omega}_{j,k}^{(t)}),\text{ and }\widehat{\Lambda}_{j,k}^{(\cdot,t)}=T_{2,\lambda_{2,j,k}^{(t)}}(\widebar{\Omega}_{j,k}^{(\cdot,t)}-\widebar{\Omega}_{j,k}^{(t)}1_{M}),$
and $\widebar{\Omega}^{(t)}=\frac{1}{N}\sum_{m=1}^{M}n_{m}\widebar{\Omega}^{(m,t)}$. Note that, the entry-adaptive shrinkage levels $\lambda_{1,j,k}^{(t)},\lambda_{2,j,k}^{(t)}\geq 0$ at the $t$-th iteration depend on the splitting proportions and the variance estimation (9), and will be specified in (26)-(27) of the subsequent section.
During each iteration $t\geq 2$, $\widetilde{\Omega}^{(m,t)}$ is sent back to the $m$-th local site for $m\in[M]$. After $T$ rounds of communication, we obtain the final integrative estimation $\widetilde{\Omega}^{(\cdot,T)}$.
Remark 3.
The data splitting strategy is employed in the debiasing steps (24) of IteHEAT. Despite the efficiency loss induced by splitting, the independence structure between $\widehat{\Omega}^{(m_{1})}$ and $\widehat{\Sigma}^{(m_{2})}$ facilitates the technical analysis, and similar approaches are utilized in Cai et al., (2023); Guo, (2023) for statistical inference. Numerically, the simulation studies in Section 6 and Section LABEL:sec:splitting_effect of the supplement demonstrate the superior performance of IteHEAT without splitting. This finding is consistent with Cai et al., (2023); Guo, (2023) and suggests that non-split IteHEAT is preferable for practical implementation.
5.2 Theoretical Properties
This section establishes the theoretical properties of the IteHEAT approach in Algorithm 5. We begin with the following additional technical assumption.
Assumption 5.
$\max_{m\in[M]}\sigma_{\max}\left(\Sigma^{(m)}\right)\leq C_{\max}$, where $\sigma_{\max}\left(\Sigma^{(m)}\right)$ is the maximal eigenvalue of $\Sigma^{(m)}$, and $X_{i}^{(m)}-\mu^{(m)}$ is sub-Gaussian with bounded sub-Gaussian norm uniformly for $m\in[M]$.
Assumption 5 is common and mild in high-dimensional statistics literature, though it is slightly stricter than Assumptions 1-2. The requirement for bounded maximal eigenvalues is assumed for simplicity and clarity, and it can be relaxed to bounded sparsely restricted eigenvalues.
Next, we establish the error analysis for the IteHEAT estimators $\widetilde{\Omega}^{(\cdot,t)},t\geq 2$ under loss matrices (17)-(18). The shrinkage levels $\lambda_{1,j,k}^{(t)},\lambda_{2,j,k}^{(t)},t\geq 2$ in (25) are specified as
$\lambda_{1,j,k}^{(t)}=(2+\delta)\sqrt{A_{j,k}\frac{\log(p)}{N}}+\widetilde{C}_{1}\left(\frac{M\log(p)}{N}+s_{0}\sqrt{\frac{\log(p)}{n}}(\max_{j,k\in[p]}\lambda_{1,j,k}^{(t-1)}+\max_{j,k\in[p]}\lambda_{2,j,k}^{(t-1)})\right),$ (26)
$\lambda_{2,j,k}^{(t)}=$ $(1+\delta)\sqrt{\frac{B_{1,j,k}+2\sqrt{2}B_{2,j,k}\sqrt{\log(p)}+4B_{\infty,j,k}\log(p)}{N}}$ (27)
$+\widetilde{C}_{2}\left(\frac{\sqrt{M}\log(p)}{\sqrt{nN}}+s_{0}\sqrt{\frac{\log(p)}{n}}(\max_{j,k\in[p]}\lambda_{1,j,k}^{(t-1)}+\max_{j,k\in[p]}\lambda_{2,j,k}^{(t-1)})\right),$
with $\lambda_{1,j,k}^{(1)}:=\lambda_{1,j,k}$ and $\lambda_{2,j,k}^{(1)}:=\lambda_{2,j,k}$, where $A_{j,k}=\sum_{m\in[M]}\frac{n_{m}}{\kappa_{m}N}\widehat{v}_{j,k}^{(m)}$, $B_{1,j,k}=\sum_{m=1}^{M}\widehat{v}_{j,k}^{(m)}/\kappa_{m}$, $B_{2,j,k}=\sqrt{\sum_{m=1}^{M}(\widehat{v}_{j,k}^{(m)}/\kappa_{m})^{2}}$, $B_{\infty,j,k}=\max_{m\in[M]}\widehat{v}_{j,k}^{(m)}/\kappa_{m}$, $\delta>0$ is given in Equations (19)-(20), and $\widetilde{C}_{1},\widetilde{C}_{2}>0$ are some universal constants independent of $(n,p,t)$. The following theorems present iterative contraction rates of convergence for IteHEAT.
Theorem 5.
Under Assumptions 1-5, if $n\geq Cs_{0}^{2}\log(p)$ for some constant $C>0$, $s_{0}\log(p)\log^{2}(p\vee n_{\max})=o(n)$, $M^{4}\log^{4}(p\vee n_{\max})=o(n)$ and $\lambda_{j}^{(m)}\asymp\sqrt{\log(p)/n_{m}}$ uniformly for $j\in[p]$ and $m\in[M]$, then for $t\geq 2$, $r\geq 1$ and $l\in[1,\infty]$, we have
$\|L_{1,r}(\widetilde{\Omega}^{(\cdot,t)},\Omega^{(\cdot)})\|_{l}\lesssim_{\mathsf{P}}s_{1}(\varphi_{1}^{(t)})^{r}+s_{2}(\varphi_{2}^{(t)})^{r},$
$\|L_{2,r}(\widetilde{\Omega}^{(\cdot,t)},\Omega^{(\cdot)})\|_{l}\lesssim_{\mathsf{P}}s_{1}(\varphi_{1}^{(t)})^{r}+s_{2}(\varphi_{2}^{(t)})^{r},$
where
$\varphi_{1}^{(t)}=\varphi_{1}^{\star}+s_{0}\sqrt{\frac{\log(p)}{n}}\varphi_{2}^{\star}+\left(s_{0}\sqrt{\frac{\log(p)}{n}}\right)^{t-1}(\varphi_{1}+\varphi_{2}),$
$\varphi_{2}^{(t)}=\varphi_{2}^{\star}+\left(s_{0}\sqrt{\frac{\log(p)}{n}}\right)^{t-1}(\varphi_{1}+\varphi_{2}),$
and $\varphi_{1}^{\star}=\sqrt{\frac{\log(p)}{N}}+\frac{M\log(p)}{N}$, $\varphi_{2}^{\star}=\sqrt{\frac{M+\log(p)}{N}}+\frac{\sqrt{M}\log(p)}{\sqrt{nN}}$.
Theorem 6.
Under the conditions in Theorem 5, we have
$\frac{1}{p}\|L_{1,r}(\widetilde{\Omega}^{(\cdot,t)},\Omega^{(\cdot)})\|_{\text{F}}^{2}\lesssim_{\mathsf{P}}s_{1}(\varphi_{1}^{(t)})^{2r}+s_{2}(\varphi_{2}^{(t)})^{2r},$
$\frac{1}{p}\|L_{2,r}(\widetilde{\Omega}^{(\cdot,t)},\Omega^{(\cdot)})\|_{\text{F}}^{2}\lesssim_{\mathsf{P}}s_{1}(\varphi_{1}^{(t)})^{2r}+s_{2}(\varphi_{2}^{(t)})^{2r}.$
The results of IteHEAT are analogous to those of HEAT in Theorems 1-2, with $\varphi_{1}$ and $\varphi_{2}$ replaced by $\varphi_{1}^{(t)}$ and $\varphi_{2}^{(t)}$, respectively. Both $\varphi_{1}^{(t)}$ and $\varphi_{2}^{(t)}$ consist of iteration-independent and iteration-dependent terms, which respectively represent the leading errors and the iteratively refined higher-order errors. Notably, the additional terms $\frac{M\log(p)}{N}$ and $\frac{\sqrt{M}\log(p)}{\sqrt{nN}}$ in $\varphi_{1}^{\star}$ and $\varphi_{2}^{\star}$ are induced by the mean estimation for $\mu^{(m)},m\in[M]$. They are included here for generality and will vanish if the means are known (e.g., $\mu^{(m)}=0$) or if the means are assumed to be homogeneous across distributed sites and are estimated using the full datasets.
The factor $s_{0}\sqrt{\log(p)/n}$ in the iteration-dependent terms serves as the contraction rate for higher-order error reduction and leads to the geometric convergence. As a result, the number of iterations $T$ with the logarithm complexity
$T\gtrsim 1+\log\left(\frac{\sqrt{M+\log(p)}+s_{0}\sqrt{M/n}\log(p)}{\sqrt{\log(p)}+M/\sqrt{N}\log(p)}\right)/\log\left(\frac{1}{s_{0}}\sqrt{\frac{n}{\log(p)}}\right)$
suffices to yield the following rates for $\widetilde{\Omega}^{(\cdot,T)}$:
$\varphi_{1}^{(T)}\asymp\varphi_{1}^{\star}+s_{0}\sqrt{\frac{\log(p)}{n}}\varphi_{2}^{\star},\text{ and }\varphi_{2}^{(T)}\asymp\varphi_{2}^{\star}.$
That is, the higher-order errors induced by aggregating local estimators are dominated by the leading errors after logarithm rounds of communication between the central and local sites. Note that, the contraction results indicate that the local sample sizes cannot be excessively small, a requirement also noted in iterative distributed algorithms under homogeneity (e.g., Wang et al.,, 2017; Jordan et al.,, 2019; Chen et al.,, 2019). Consequently, with a few rounds of iterations, the results in Theorems 5-6 match the minimax lower bounds in Theorems 3-4 if $\frac{M^{2}}{N}\lesssim(\log(p))^{-1}$, $\log^{2}(p)\lesssim(1\vee\frac{\log(p)}{M})n$ and $M\vee\log(p)\lesssim ns_{0}^{-2}$. Therefore, IteHEAT enhances accuracy and optimality compared to HEAT by leveraging iterative refinements.
6 Simulation Study
[…]
Note that HEAT estimation is essentially the first-round output of IteHEAT. Therefore, we present the performance of the IteHEAT estimators $\widetilde{\Omega}^{(\cdot,t)}$ for $t=0,1,2,\ldots$, where $\widetilde{\Omega}^{(\cdot,1)}$ represents the HEAT estimator, and $\widetilde{\Omega}^{(\cdot,0)}:=\widehat{\Omega}^{(\cdot)}$ represents the individual local estimator in (4)-(5). The practical determination of regularization parameters $\lambda_{j}^{(m)},j\in[p],m\in[M]$ in Lasso procedures (3) and (22), as well as the shrinkage levels $\lambda_{1,j,k},\lambda_{2,j,k}$ and $\lambda_{1,j,k}^{(t)},\lambda_{2,j,k}^{(t)},t=2,3,\ldots$ in HEAT (12) and IteHEAT (25), is discussed in Sections LABEL:sec:reg_para-LABEL:sec:shrink_para of the supplement. Section LABEL:sec:splitting_effect of the supplement evaluates the effects of data splitting on the proposed IteHEAT and shows that IteHEAT without splitting exhibits superior performance. Therefore, throughout this section, IteHEAT is implemented without data splitting. We employ the soft and SCAD thresholding functions, and denote the corresponding methods by HEAT & IteHEAT_Soft and HEAT & IteHEAT_SCAD, respectively. For illustration purpose, we set $n_{0}=400,p\in\{100,200\}$ and $M\in\{5,20\}$.
[…]
7 Real Data Analysis
In this section, we explore the version 8 bulk tissue RNA-seq count data in Genotype-Tissue Expression (GTEx) project. This dataset collects samples from 54 tissues across nearly 1000 donors, and can be accessed publicly at https://www.gtexportal.org. We focus on nine tissues with the most frequently collected samples: Adipose-Subcutaneous, Lung, Muscle-Skeletal, Skin-Sun Exposed, Skin-Not Sun Exposed, Thyroid, Nerve-Tibial, Artery-Tibial and Whole Blood (GTEx-Consortium,, 2020), and then employ our proposed methodology to conduct differential analysis for expression networks. Gender differences in immune diseases are widely acknowledged (Moroni et al.,, 2012; Ngo et al.,, 2014). Thus, we investigate gender heterogeneity within each tissue separately, focusing on the gene set related to antigen processing and presentation (GO:0019882), which contains 117 genes. Consequently, we have $M=2$ across tissues. The sample sizes in male datasets range from 395 to 543, while those in female datasets range from 183 to 260.
Figure 3: Top ten genes showing highest heterogeneous connectivity in each tissue. The y-axis shows the number of heterogeneous edges. The variables n_F, n_M, and p represent the numbers of female samples, male samples and remaining genes in each tissue, respectively.
For each tissue analysis, we employ edgeR (Robinson et al.,, 2010) to preprocess the RNA-seq count datasets. This includes low-expressed gene filtering, normalization, and log-counts-per-million transformation. After preprocessing, the numbers of remaining genes range from 100 to 111. Due to the limited number of integrated datasets ($M=2$), we implement HEAT with SCAD thresholding separately for each tissue. We then extract estimated network edges that exhibit gender heterogeneity. The top ten genes showing the highest connectivity within heterogeneous edges are illustrated in Figure 3. The results reveal that the gene KDM5D exhibits the highest degree of heterogeneity in connectivity. Additionally, heterogeneous associations related to the HLA gene family are observed in several tissues. Notably, KDM5D, located on the Y chromosome, is recognized as a tumor suppressor (Ohguchi and Ohguchi,, 2022), and HLA genes are associated with various immune diseases (Ngo et al.,, 2014). These findings of gender-related genetic association heterogeneity support the validity of the proposed methodology.
8 Conclusion and Discussion
In this article, we focus on distributed integrative estimation for high-dimensional precision matrices under heterogeneity and introduce the HEAT and IteHEAT methods. These proposals achieve both communication and computation efficiency while maintaining the statistical optimality. Importantly, the ideas underlying HEAT and IteHEAT are not limited to precision matrix estimation alone, and can be extended to various models with unbiased estimators that can be facilitated by modern inference theories (e.g., Chernozhukov et al.,, 2018).
In conclusion, we raise two questions for future research. First, while simulations demonstrate the superiority of IteHEAT without data splitting, a rigorous theoretical guarantee is beneficial to validate its performance. Second, the current heterogeneity modelling in HEAT and IteHEAT may be sensitive to highly adversarial or contaminated datasets. Therefore, developing robust methodologies for distributed integration warrants further exploration.